# HiRAG: Retrieval-Augmented Generation with Hierarchical Knowledge

### v2

@@ Abstract
Graph-based Retrieval-Augmented Generation (RAG) methods have significantly enhanced the performance of large language models (LLMs) in domain-specific tasks. However, existing RAG methods do not adequately utilize the naturally inherent hierarchical knowledge in human cognition, which limits the capabilities of RAG systems. In this paper, we introduce a new RAG approach, called HiRAG, which utilizes hierarchical knowledge to enhance the semantic understanding and structure capturing capabilities of RAG systems in the indexing and retrieval processes. Our extensive experiments demonstrate that HiRAG achieves significant performance improvements over the state-of-the-art baseline methods.111 https://github.com/hhy-huang/HiRAG
@@ 1 Introduction
Retrieval Augmented Generation (RAG) Gao et al. (2023) Lewis et al. (2020) Fan et al. (2024) has been introduced to enhance the capabilities of LLMs in domain-specific or knowledge-intensive tasks. Naive RAG methods retrieve text chunks that are relevant to a query, which serve as references for LLMs to generate responses, thus helping address the problem of "Hallucination" Zhang et al. (2023) Tang and Yang (2024). However, naive RAG methods usually overlook the relationships among entities in the retrieved text chunks. To address this issue, RAG systems with graph structures were proposed Edge et al. (2024) Liang et al. (2024) Zhang et al. (2025) Peng et al. (2024b), which construct knowledge graphs (KGs) to model relationships between entities in the input documents. Although existing RAG systems integrating graph structures have demonstrated outstanding performance on various tasks, they still have some serious limitations. GraphRAG Edge et al. (2024) introduces communities in indexing using the Leiden algorithm Traag et al. (2019), but the communities only capture the structural proximity of the entities in the KG. KAG Liang et al. (2024) indexes with a hierarchical representation of information and knowledge, but their hierarchical structure relies too much on manual annotation and requires a lot of human domain knowledge, which renders their method not scalable to general tasks. LightRAG Guo et al. (2024) utilizes a dual-level retrieval approach to obtain local and global knowledge as the contexts for a query, but it ignores the knowledge gap between local and global knowledge, that is, local knowledge represented by the retrieved individual entities (i.e., entity-specific details) may not be semantically related to the global knowledge represented in the retrieved community summaries (i.e., broader, aggregated summaries), as these individual entities may not be a part of the retrieved communities for a query.
@@
Challenge (2) occurs as existing methods Guo et al. (2024) Edge et al. (2024) typically retrieve context either from global or local perspectives but fail to address the inherent disparity between these knowledge layers. Consider the query "Please introduce Amazon" in Figure 1, where global context emphasizes Amazon’s involvement in technological domains like big data and cloud computing, but local context retrieves entities directly linked to Amazon (e.g., subsidiaries, leadership). When these two knowledge layers are fed into LLMs as the contexts of a query without contextual alignment, LLMs may struggle to reconcile their distinct scopes, leading to disjointed reasoning, incomplete answers, or even contradictory outputs. For instance, an LLM might conflate Amazon’s role as a cloud provider (global) with its e-commerce operations (local), resulting in incoherent or factually inconsistent responses as the red words shown in the case. This underscores the need for new methods that bridge hierarchical knowledge layers to ensure cohesive reasoning in RAG systems.
@@ 2 Related Work
In this section, we discuss recent research concerning graph-augmented LLMs, specifically RAG methods with graph structures. GNN-RAG Mavromatis and Karypis (2024) employs GNN-based reasoning to retrieve query-related entities. Then they find the shortest path between the retrieved entities and candidate answer entities to construct reasoning paths. LightRAG Guo et al. (2024) integrates a dual-level retrieval method with graph-enhanced text indexing. They also decrease the computational costs and speed up the adjustment process. GRAG Hu et al. (2024) leverages a soft pruning approach to minimize the influence of irrelevant entities in retrieved subgraphs. It also implements prompt tuning to help LLMs comprehend textual and topological information in subgraphs by incorporating graph soft prompts. StructRAG Li et al. (2024) identifies the most suitable structure for each task, transforms the initial documents into this organized structure, and subsequently generates responses according to the established structure. Microsoft GraphRAG Edge et al. (2024) first retrieves related communities and then let the LLM generate the response with the retrieved communities. They also answer a query with global search and local search. KAG Liang et al. (2024) introduces a professional domain knowledge service framework and employs knowledge alignment using conceptual semantic reasoning to mitigate the noise issue in OpenIE. KAG also constructs domain expert knowledge using human-annotated schemas. HippoRAG Gutiérrez et al. (2024) synergistically integrates LLMs, KGs, and Personalized PageRank to mimic the neocortex and hippocampus roles in human memory.
@@ 3 Preliminary and Definitions
In this section, we present the formulation of a graph-based RAG system, extending the definitions established in Guo et al. (2024) and Peng et al. (2024a).
In a graph-based RAG system framework as shown in Equation 1, is the generation module, represents the retrieval module, denotes the graph indexer, and refers to the graph retriever:
When we answer a query , the graph-based RAG system produces an optimal response by maximizing a target distribution that captures the probability of generating each candidate response. Formally, this is defined as:
where is the target distribution with a graph retriever and a generator , and is a set of possible responses. The graph database is constructed from the original external database . We utilize the total probability formula to decompose , which can be expressed as
@@ 4.2 Retrieval with Hierarchical Knowledge
After retrieving the three-level hierarchical knowledge, i.e., local-level descriptions of the individual entities in , global-level community reports of the communities in , and bridge-level descriptions of the triples in , we feed them as the context to the LLM to generate a comprehensive answer to the query. We also provide the detailed procedures of HiRAG with pseudocodes in Appendix D.
@@ 5 Experimental Evaluation
In this section. We report the performance evaluation results of HiRAG on Query-Focused Summarization (QFS) and Multi-Hop QA (MHQA) tasks respectively.
Baseline Methods. We compared HiRAG with state-of-the-art and popular baseline RAG methods, which are designed for QFS and MHQA tasks respectively. For QFS tasks: NaiveRAG Gao et al. (2022) Gao et al. (2023) splits original documents into chunks and retrieves relevant text chunks through vector search. GraphRAG Edge et al. (2024) utilizes communities and we use the local search mode in our experiments as it retrieves community reports as global knowledge, while their global search mode is known to be too costly and does not use local entity descriptions. LightRAG Guo et al. (2024) uses both global and local knowledge to answer a query. FastGraphRAG Circlemind (2024) integrates KG and personalized PageRank as proposed in HippoRAG Gutiérrez et al. (2024). KAG Liang et al. (2024) integrates structured reasoning of KG with LLMs and employs mutual indexing and logical-form-guided reasoning to enhance professional domain knowledge services. For MHQA tasks: in addition to NaiveRAG, GraphRAG, LightRAG, and FastGraphRAG, we also include RAPTOR, HippoRAG Gutiérrez et al. (2024) and HippoRAG2 Gutiérrez et al. (2025) as the state-of-the-art RAG methods for MHQA tasks.
Datasets and Queries. For QFS tasks, we used four datasets from the UltraDomain benchmark Qian et al. (2024), which is designed to evaluate RAG systems across diverse applications, focusing on long-context tasks and high-level queries in specialized domains. We used Mix, CS, Legal, and Agriculture datasets like in LightRAG Guo et al. (2024). We also used the benchmark queries provided in UltraDomain for each of the four datasets. For MHQA tasks, we randomly sampled 1,000 queries from 2WikiMultiHopQA Ho et al. (2020) and HotpotQA Yang et al. (2018) following the settings in HippoRAG2 Gutiérrez et al. (2025). The statistics of these datasets are given in Appendix B.
LLM. For QFS tasks, we employed DeepSeek-V3 DeepSeek-AI et al. (2024) as the LLM for information extraction, entity summarization, and answer generation in HiRAG and other baseline methods. We utilized GLM-4-Plus GLM et al. (2024) as the embedding model for vector search and semantic clustering because DeepSeek-V3 does not provide an accessible embedding model. For MHQA tasks, we used GPT-4o-mini and nvidia/NVEmbed-v2 Lee et al. (2024) as the LLM and embedding model used in HiRAG and other baseline methods respectively, following the same settings in HippoRAG2.
@@
| 2WikiMultiHopQA | HotpotQA | Average | ||||
| Method | EM | F1 | EM | F1 | EM | F1 |
| NaiveRAG | 54.4 | 60.8 | 57.3 | 71.0 | 55.9 | 65.9 |
| RAPTOR | 39.7 | 48.4 | 50.6 | 64.7 | 45.2 | 56.6 |
| GraphRAG | 45.7 | 61.0 | 51.4 | 67.6 | 48.6 | 64.3 |
| LightRAG | 2.5 | 12.1 | 9.9 | 20.2 | 6.3 | 16.2 |
| FastGraphRAG | 20.8 | 44.8 | 35.0 | 49.6 | 27.9 | 47.2 |
| HippoRAG | 59.4 | 67.3 | 46.3 | 60.0 | 52.9 | 63.7 |
| HippoRAG2 | 60.5 | 69.7 | 56.3 | 71.1 | 58.4 | 70.4 |
| HiRAG | 69.0 | 74.4 | 62.0 | 72.9 | 65.5 | 73.7 |
@@ 5.1 Overall Performance Comparison
QFS Task Evaluation Details. For QFS tasks, our experiments followed the evaluation methods of recent work Edge et al. (2024)Guo et al. (2024) by employing a powerful LLM to conduct multi-dimensional comparison. We used the win rate to compare different methods, which indicates the percentage of instances that a method generates higher-quality answers compared to another method as judged by the LLM. We utilized GPT-4o Achiam et al. (2023) as the evaluation model to judge which method generates a superior answer for each query for the following four dimensions: (1) Comprehensiveness: how thoroughly does the answer address the question, covering all relevant aspects and details? (2) Empowerment: how effectively does the answer provide actionable insights or solutions that empower the user to take meaningful steps? (3) Diversity: how well does the answer incorporate a variety of perspectives, approaches, or solutions to the problem? (4) Overall: how does the answer perform overall, considering comprehensiveness, empowerment, diversity, and any other relevant factors? For a fair comparison, we also alternated the order of the answers generated by each pair of methods in the prompts and calculated the overall win rates of each method. We also report the reasonability and credibility of using a powerful LLM as a judge in Appendix G.
MHQA Task Evaluation Details. For MHQA tasks, our experiments also followed the evaluation methods of recent work Gutiérrez et al. (2024) Gutiérrez et al. (2025) by leveraging two established metrics: Exact Match (EM) and F1 scores, which are applied to the generated answers. Compared with the metric of win rates used in QFS tasks, the performance with EM and F1 scores can indicate HiRAG’s ability to achieve correctness.
Evaluation Results. We present the win rates, EM, and F1 scores of HiRAG and various baseline methods in Table 1 and Table 2. HiRAG consistently outperforms existing approaches in both QFS tasks across all four datasets and four evaluation dimensions and MHQA tasks across all two datasets in the majority of cases. Key insights derived from the results are summarized below.
@@
To show the effectiveness of HiRetrieval, we also created another variant of HiRAG without using the bridge-level knowledge, denoted by w/o Bridge in Table 1. The result shows that without the bridge-layer knowledge, the win rates drop significantly across all datasets and evaluation dimensions, because there is knowledge gap between the local-level and global-level knowledge as discussed in Section 1. We also report the knowledge coverage of bridge-level descriptions in Appendix H, which further proofs both local- and global-level knowledge are well connected in the bridge-level descriptions.
@@ 5.4 Determining the Number of Layers
5.5 Efficiency and Costs Analysis
To evaluate the efficiency and costs of HiRAG, we also report the token costs, the number of API calls, and the time costs of indexing and retrieval of HiRAG and the baselines in Table 5 in Appendix C. Although HiRAG needs more time and resources to conduct indexing for better performance, we remark that indexing is offline and the total cost is only about 7.55 USD for the Mix dataset using DeepSeek-V3. In terms of retrieval, unlike KAG and LightRAG, HiRAG does not cost any tokens for retrieval. Therefore, HiRAG is more efficient for online retrieval.
6 Conclusions
Appendix
Appendix A Limitations
HiRAG has the following limitations. Firstly, constructing a high-quality hierarchical KG may incur substantial token consumption and time overhead, as LLMs need to perform entity summarization in each layer. However, the monetary cost of using LLMs may not be the major concern as the cost is decreasing rapidly recently, and therefore we may consider parallelizing the indexing process to reduce the indexing time. Secondly, the retrieval module requires more sophisticated query-aware ranking mechanisms. Currently, our HiRetrieval module relies solely on LLM-generated weights for relation ranking, which may affect query relevance. We will research for more effective ranking mechanisms to further improve the retrieval quality.
Appendix B Experimental Datasets
| Dataset | Mix | CS | Legal | Agriculture |
| # of Documents | 61 | 10 | 94 | 12 |
| # of Tokens | 625,948 | 2,210,894 | 5279400 | 2,028,496 |
| Dataset | 2WikiMultiHopQA | HotpotQA |
| # of Queries | 1,000 | 1,000 |
| # of Passages | 6,119 | 9,811 |
| # of Tokens | 1,189,866 | 1,507,487 |
Appendix C Cost Analysis
@@
As shown in Table 5, we calculate the average token, API calls and time costs of different methods across four datasets. For indexing, we record the total costs of the entire indexing process. For retrieval, we calculate the average costs per query during the retrieval process, which could reflect the performance while the methods are deployed online.
Appendix D Implementation Details of HiRAG

### v3

# Retrieval-Augmented Generation with Hierarchical Knowledge

@@ Abstract
Graph-based Retrieval-Augmented Generation (RAG) methods have significantly enhanced the performance of large language models (LLMs) in domain-specific tasks. However, existing RAG methods do not adequately utilize the naturally inherent hierarchical knowledge in human cognition, which limits the capabilities of RAG systems. In this paper, we introduce a new RAG approach, called HiRAG, which utilizes hierarchical knowledge to enhance the semantic understanding and structure capturing capabilities of RAG systems in the indexing and retrieval processes. Our extensive experiments demonstrate that HiRAG achieves significant performance improvements over the state-of-the-art baseline methods. The code of our proposed method is available at https://github.com/hhy-huang/HiRAG.
Retrieval-Augmented Generation with Hierarchical Knowledge
Haoyu Huang1,2111Equal contribution. This research was conducted at kasma.ai., Yongfeng Huang2111Equal contribution. This research was conducted at kasma.ai., Junjie Yang2, Zhenyu Pan1,2, Yongqiang Chen1 Kaili Ma1, Hongzhi Chen1, James Cheng2 1KASMA.ai 2CSE, The Chinese University of Hong Kong {haoyuhuang,zhenyupan,yqchen,klma,chenhongzhi}@kasma.ai {haoyuhuang,zhenyupan,1155215805}@link.cuhk.edu.hk {yfhuang22,jcheng}@cse.cuhk.edu.hk
@@ 1 Introduction
Retrieval Augmented Generation (RAG) Gao et al. (2023); Lewis et al. (2020); Fan et al. (2024) has been introduced to enhance the capabilities of LLMs in domain-specific or knowledge-intensive tasks. Naive RAG methods retrieve text chunks that are relevant to a query, which serve as references for LLMs to generate responses, thus helping address the problem of "Hallucination" Zhang et al. (2023); Tang and Yang (2024). However, naive RAG methods usually overlook the relationships among entities in the retrieved text chunks. To address this issue, RAG systems with graph structures were proposed Edge et al. (2024); Liang et al. (2024); Zhang et al. (2025); Peng et al. (2024a), which construct knowledge graphs (KGs) to model relationships between entities in the input documents. Although existing RAG systems integrating graph structures have demonstrated outstanding performance on various tasks, they still have some serious limitations. GraphRAG Edge et al. (2024) introduces communities in indexing using the Leiden algorithm Traag et al. (2019), but the communities only capture the structural proximity of the entities in the KG. KAG Liang et al. (2024) indexes with a hierarchical representation of information and knowledge, but their hierarchical structure relies too much on manual annotation and requires a lot of human domain knowledge, which renders their method not scalable to general tasks. LightRAG Guo et al. (2024) utilizes a dual-level retrieval approach to obtain local and global knowledge as the contexts for a query, but it ignores the knowledge gap between local and global knowledge, that is, local knowledge represented by the retrieved individual entities (i.e., entity-specific details) may not be semantically related to the global knowledge represented in the retrieved community summaries (i.e., broader, aggregated summaries), as these individual entities may not be a part of the retrieved communities for a query.
@@
Challenge (2) occurs as existing methods Guo et al. (2024); Edge et al. (2024) typically retrieve context either from global or local perspectives but fail to address the inherent disparity between these knowledge layers. Consider the query "Please introduce Amazon" in Figure 1, where global context emphasizes Amazon’s involvement in technological domains like big data and cloud computing, but local context retrieves entities directly linked to Amazon (e.g., subsidiaries, leadership). When these two knowledge layers are fed into LLMs as the contexts of a query without contextual alignment, LLMs may struggle to reconcile their distinct scopes, leading to disjointed reasoning, incomplete answers, or even contradictory outputs. For instance, an LLM might conflate Amazon’s role as a cloud provider (global) with its e-commerce operations (local), resulting in incoherent or factually inconsistent responses as the red words shown in the case. This underscores the need for new methods that bridge hierarchical knowledge layers to ensure cohesive reasoning in RAG systems.
@@ 2 Related Work
In this section, we discuss recent research concerning graph-augmented LLMs, specifically RAG methods with graph structures. GNN-RAG Mavromatis and Karypis (2024) employs GNN-based reasoning to retrieve query-related entities. Then they find the shortest path between the retrieved entities and candidate answer entities to construct reasoning paths. LightRAG Guo et al. (2024) integrates a dual-level retrieval method with graph-enhanced text indexing. They also decrease the computational costs and speed up the adjustment process. GRAG Hu et al. (2024) leverages a soft pruning approach to minimize the influence of irrelevant entities in retrieved subgraphs. It also implements prompt tuning to help LLMs comprehend textual and topological information in subgraphs by incorporating graph soft prompts. StructRAG Li et al. (2024) identifies the most suitable structure for each task, transforms the initial documents into this organized structure, and subsequently generates responses according to the established structure. Microsoft GraphRAG Edge et al. (2024) first retrieves related communities and then let the LLM generate the response with the retrieved communities. They also answer a query with global search and local search. KAG Liang et al. (2024) introduces a professional domain knowledge service framework and employs knowledge alignment using conceptual semantic reasoning to mitigate the noise issue in OpenIE. KAG also constructs domain expert knowledge using human-annotated schemas. ReG Zou et al. (2025) uses LLMs to refine weak supervision signals to align retrievers in fine-grained graph-based RAG systems.
@@ 3 Preliminary and Definitions
In this section, we give a general formulation of an RAG system with graph structure referring to the definitions in Guo et al. (2024) and Peng et al. (2024b).
In an RAG framework as shown in Equation 1, LLM is the generation module, represents the retrieval module, denotes the graph indexer, and refers to the graph retriever:
When we answer a query, the answer we get from an RAG system is represented by , which can be formulated as
where is the target distribution with a graph retriever and a generator , and is the set of possible responses. The graph database is constructed from the original external database . We utilize the total probability formula to decompose , which can be expressed as
@@ 4.2 Retrieval with Hierarchical Knowledge
After retrieving the three-level hierarchical knowledge, i.e., local-level descriptions of the individual entities in , global-level community reports of the communities in , and bridge-level descriptions of the triples in , we feed them as the context to the LLM to generate a comprehensive answer to the query. We also provide the detailed procedures of HiRAG with pseudocodes in Appendix C.
@@ 5 Experimental Evaluation
We report the performance evaluation results of HiRAG in this section.
Baseline Methods. We compared HiRAG with state-of-the-art and popular baseline RAG methods. NaiveRAG Gao et al. (2022, 2023) splits original documents into chunks and retrieves relevant text chunks through vector search. GraphRAG Edge et al. (2024) utilizes communities and we use the local search mode in our experiments as it retrieves community reports as global knowledge, while their global search mode is known to be too costly and does not use local entity descriptions. LightRAG Guo et al. (2024) uses both global and local knowledge to answer a query. FastGraphRAG Circlemind (2024) integrates KG and personalized PageRank as proposed in HippoRAG Gutiérrez et al. (2024). KAG Liang et al. (2024) integrates structured reasoning of KG with LLMs and employs mutual indexing and logical-form-guided reasoning to enhance professional domain knowledge services.
Datasets and Queries. We used four datasets from the UltraDomain benchmark Qian et al. (2024), which is designed to evaluate RAG systems across diverse applications, focusing on long-context tasks and high-level queries in specialized domains. We used Mix, CS, Legal, and Agriculture datasets like in LightRAG Guo et al. (2024). We also used the benchmark queries provided in UltraDomain for each of the four datasets. The statistics of these datasets are given in Appendix A.
LLM. We employed DeepSeek-V3 DeepSeek-AI et al. (2024) as the LLM for information extraction, entity summarization, and answer generation in HiRAG and other baseline methods. We utilized GLM-4-Plus GLM et al. (2024) as the embedding model for vector search and semantic clustering because DeepSeek-V3 does not provide an accessible embedding model.
@@ 5.1 Overall Performance Comparison
Evaluation Details. Our experiments followed the evaluation methods of recent work Edge et al. (2024); Guo et al. (2024) by employing a powerful LLM to conduct multi-dimensional comparison. We used the win rate to compare different methods, which indicates the percentage of instances that a method generates higher-quality answers compared to another method as judged by the LLM. We utilized GPT-4o Achiam et al. (2023) as the evaluation model to judge which method generates a superior answer for each query for the following four dimensions: (1) Comprehensiveness: how thoroughly does the answer address the question, covering all relevant aspects and details? (2) Empowerment: how effectively does the answer provide actionable insights or solutions that empower the user to take meaningful steps? (3) Diversity: how well does the answer incorporate a variety of perspectives, approaches, or solutions to the problem? (4) Overall: how does the answer perform overall, considering comprehensiveness, empowerment, diversity, and any other relevant factors? For a fair comparison, we also alternated the order of the answers generated by each pair of methods in the prompts and calculated the overall win rates of each method.
Evaluation Results. We present the win rates of HiRAG and five baseline methods in Table 1. HiRAG consistently outperforms existing approaches across all four datasets and four evaluation dimensions in the majority of cases. Key insights derived from the results are summarized below.
@@
| Mix | CS | Legal | Agriculture | |
| Recall (%) | Recall (%) | Recall (%) | Recall (%) | |
| Global | 38.61 | 48.96 | 53.44 | 50.75 |
| Local | 83.07 | 81.88 | 78.13 | 64.47 |
@@
To show the effectiveness of HiRetrieval, we also created another variant of HiRAG without using the bridge-level knowledge, denoted by w/o Bridge in Table 1. The result shows that without the bridge-layer knowledge, the win rates drop significantly across all datasets and evaluation dimensions, because there is knowledge gap between the local-level and global-level knowledge as discussed in Section 1. We also report the knowledge coverage of bridge-level descriptions over global- and local-level knowledge in Table 2 and Appendix G, which further proofs both local- and global-level knowledge are well connected in the bridge-level descriptions.
@@ 5.4 Determining the Number of Layers
5.5 Efficiency and Costs Analysis
To evaluate the efficiency and costs of HiRAG, we also report the token costs, the number of API calls, and the time costs of indexing and retrieval of HiRAG and the baselines in Table 3. For indexing, we record the total costs of the entire indexing process. For retrieval, we calculate the average costs per query during the retrieval process, which could reflect the performance while the methods are deployed online.
Although HiRAG needs more time and resources to conduct indexing for better performance, we remark that indexing is offline and the total cost is only about 7.55 USD for the Mix dataset using DeepSeek-V3. In terms of retrieval, unlike KAG and LightRAG, HiRAG does not cost any tokens for retrieval. Therefore, HiRAG is more efficient for online retrieval.
6 Conclusions
7 Limitations
HiRAG has the following limitations. Firstly, constructing a high-quality hierarchical KG may incur substantial token consumption and time overhead, as LLMs need to perform entity summarization in each layer. However, the monetary cost of using LLMs may not be the major concern as the cost is decreasing rapidly recently, and therefore we may consider parallelizing the indexing process to reduce the indexing time. Secondly, the retrieval module requires more sophisticated query-aware ranking mechanisms. Currently, our HiRetrieval module relies solely on LLM-generated weights for relation ranking, which may affect query relevance. In the future, we will research for more effective ranking mechanisms to further improve the retrieval quality. Besides, we can also incorporate causality into HiRAG to enhance reasoning capabilities of LLMs Liu et al. (2024); Chen et al. (2022).
Appendix
Appendix A Experimental Datasets
| Dataset | Mix | CS | Legal | Agriculture |
| # of Documents | 61 | 10 | 94 | 12 |
| # of Tokens | 625948 | 2210894 | 5279400 | 2028496 |
Table 4 presents the statistical characteristics of the experimental datasets, where all documents were consistently tokenized using Byte Pair Encoding (BPE) tokenizer "cl100k_base".
Appendix B Evaluations with Objective Metrics
| 2WikiMultiHopQA | HotpotQA | |||
| Method | EM (%) | F1 (%) | EM (%) | F1 (%) |
| NaiveRAG | 15.60 | 25.64 | 21.60 | 40.19 |
| GraphRAG | 22.50 | 27.49 | 31.70 | 42.74 |
| LightRAG | 16.50 | 40.95 | 25.00 | 43.20 |
| FastGraphRAG | 20.80 | 44.81 | 35.00 | 49.56 |
| HiRAG | 46.20 | 60.06 | 37.00 | 52.29 |
To objectively evaluate the QA performance of HiRAG and the baseline methods, we leverage two established metrics: exact match (EM) and F1 scores, applied to the generated answers. We perform systematic evaluations using GPT-4o-mini on two multi-hop QA datasets: HotpotQA Yang et al. (2018) and 2WikiMultiHopQA Ho et al. (2020). For a consistent comparison with previous work, we follow the settings of HippoRAG Gutiérrez et al. (2024), obtaining 1,000 queries from each validation set. We did not present the results of KAG because, despite our efforts to implement it, we were unable to make it fully work on this benchmark.
Compared with the metric of win rates, the performances with EM and F1 scores can indicate HiRAG’s ability to achieve correctness. Given that the RAG system has access to richer contexts, it tends to produce more comprehensive responses. Nevertheless, while comprehensiveness, empowerment, and diversity are important qualities for the generated answers, correctness is equally essential. As illustrated in Table 5, HiRAG is also capable of generating more accurate answers compared to the baseline methods.
Appendix C Implementation Details of HiRAG